A door with width 1.2 m and mass M = 10 kg is hinged on one side so that it can rotate


freely. A bullet is fired into the center of the door. The bullet has mass 25 g and a speed of


300 m/s. What is the angular velocity of the door with respect to the hinge just after the


bullets embeds itself in the door?

1
Expert's answer
2021-12-09T09:13:45-0500

By energy conservation:


"\frac12 mv^2=\frac12I\omega^2,\\\space\\
\frac12 mv^2=\frac12·\frac13ML^2·\omega^2,\\\space\\
\omega=\frac vL\sqrt{\frac{m}{M}}=12.5\text{ rad/s}."